OpenStudy (anonymous):

a company pays its employees a base salary of $1000 every other week. plus $30 for each sale. How many sales did the employee have if the paycheck is $1720

OpenStudy (anonymous):

1000+30x=1720 -1000 -1000 30x=720 x= 24